Establish the initial feasible solutions of the LP model using NCR (North west Corner Rule), MCM ( Minimum Cost Method or Greedy Method), and VAM ( Vogel’s Approximation Method)

 Minimize: C = 14X1A + 25X1B + 13X1C + 18 X1D.+ 10X2A + 12X2B + 13X2C + 11X2D+ 15X3A + 20X3B + 11X3C+ 25X3D

Subject to: X1A + X1B + X1C + X1D = 140  X1A + X2A + X3A = 100

                X2A + X2B + X2C + X2D = 160     X1B + X2B + X3B = 100

               X3A + X3B + X3C + X3D = 50      X1C + X2C + X3C = 50

              X1D + X2D + X3D = 100 Xij    



1
Expert's answer
2021-08-01T09:34:51-0400

"\text{The initial tableau for this problem is: }\\\begin{matrix}
 & &&&140 \\\hline
 &&&&160\\\hline&&&&50\\\hline100&100&50&100
\end{matrix}
\\\text{Using North-west corner method, we locate the upper-most variable$(X_{1A})$ in the }\\\text{tableau and set it to the min(100,140) = 100 and we cancel out column 1}
\\\begin{matrix}
 100& &&&40 \\\hline
 &&&&160\\\hline&&&&50\\\hline X&100&50&100
\end{matrix}
\\\text{Repeating the process, we obtain the following tableaus}
\\\begin{matrix}
 100& 40&&&X\\\hline
 &&&&160\\\hline&&&&50\\\hline X&60&50&100
\end{matrix}
\\\begin{matrix}
 100& 40&&&X\\\hline
 &&&&100\\\hline&&&&50\\\hline X&X&50&100
\end{matrix}
\\

\\\begin{matrix}
 100& 40&&&X\\\hline
 &60&50&&50\\\hline&&&&50\\\hline X&X&X&100
\end{matrix}


\\\begin{matrix}
 100& 40&&&X\\\hline
 &60&50&50&X\\\hline&&&&50\\\hline X&X&X&50
\end{matrix}

\\\begin{matrix}
 100& 40&&&X\\\hline
 &60&50&50&X\\\hline&&&50&X\\\hline X&X&X&X
\end{matrix}
\\\text{Therefore, the solution is feasible, the solution is $X_{1A}=100,X_{1B}=40,X_{2B}=60$,}\\\text{$X_{2C}=50,X_{2D}=50,X_{3D}=50,C=5570$}
\\\text{We locate the smallest cost(10) at $X_{2A}$ in our initial matrix, we then set the value of $X_{2A}$}\\\text{to the min(100,160). Next,we cross out column 1 and reduce $S_2$(160) to 160-100=60. We }\\\text{repeat the process for the remaining uncrossed out rows and columns to obtain the final tableau}\\
\\\begin{matrix}
 & 100&&40&X\\\hline
 100 &&&60&X\\\hline&0&50&&X\\\hline X&X&X&X
\end{matrix}
\\\text{Therefore, the initial feasible solution is }\\X_{1B}=100,X_{1D}=40,X_{2A}=100,X_{2D}=60, X_{3B}=0,X_{3C}=50,z=5430
\\\text{Next, we use the Vogel's approximation method to find the basic feasible solution. We }\\\text{compute the penalties by computing the difference between the smallest numbers in each}\\\text{row, we select the row with the largest penalty, after which we select the variable with }\\\text{the smallest value.}\\\text{From the table below, column 3 has the largest penalty, we choose $X_{3C}$ since has the smallest}\\\text{cost}
\\\begin{matrix}
 14 & 25&13&18&140 &&&14-13=1 \\\hline 
 10 &12&13&11&160&&&11-10=1\\\hline15&20&11&25&50&&&15-11=4\\\hline100&100&50&100&&&
\end{matrix}
\\\text{Repeating the process, we obtain the final tableau}
\\\begin{matrix}
 100 &&&40&X\\\hline
 &100&&60&X\\\hline0&&50&&X\\\hline X&X&X&X
\end{matrix}
\\X_{1A}=100,X_{1D}=40,X_{2B}=100,X_{2D}=60, X_{3A}=0,X_{3C}=50,z=4530"